# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ 9 Conclusion
In this paper, a detailed and comprehensive review of RSISR methods is presented, covering methodology, datasets, and evaluation metrics. Specifically, we carefully group the existing RSISR methods into supervised and unsupervised. Furthermore, representative methods in each category are analyzed and discussed in detail. We also discuss the recent developments in this field based on a survey of more than 300 papers. Through these reviews and analyses, we provide future perspectives on some issues we consider critical. We expect this survey to provide a suitable reference for researchers in the field of RSISR.

### v3

@@ 10 Conclusion
In this paper, a detailed and comprehensive review of RSISR methods is presented, covering methodology, datasets, and evaluation metrics. Specifically, we carefully group the existing RSISR methods into supervised and unsupervised. Furthermore, representative methods in each category are analyzed and discussed in detail. We also discuss the recent developments in this field based on a survey of more than 400 papers. Through these reviews and analyses, we provide future perspectives on some issues we consider critical. We expect this survey to provide a suitable reference for researchers in the field of RSISR.